OpenStudy (anonymous):

How do you start off this problem..? 2√(x+13)+5=17 (Only the [x+13] is under the square root.)

OpenStudy (anonymous):

isloate the [x+13] part

OpenStudy (anonymous):

then square it

OpenStudy (anonymous):

I thought you'd have to divide everything by 2..

OpenStudy (anonymous):

not needed

OpenStudy (alexwee123):

2√(x+13)+5=17 2√(x+13)=12 √(x+13)=6 x+13=6^2 x=36-13 x=23

OpenStudy (anonymous):

So, you first subtracted both sides by 5. Then, divided by 2. Squared both sides. And subtracted 13..

OpenStudy (alexwee123):

yes